for a boy having 6 friends how many ways are there to invite 1 or more of them to the lunch?
Answer by stanbon(75887)

for a boy having 6 friends how many ways are there to invite 1 or more of them to the lunch?
----
#(at least one) = 1 - #(none)
---
Number of ways to invite any number from 0 to 6 = 2^6 = 64
Number of ways to invite none of the 6: 1
------------
Number of ways to invite at least one: 64-1 = 63
====================================================
Cheers,
Stan H.